# RangeSAM: Leveraging Visual Foundation Models for Range-View repesented LiDAR segmentation

### v1

@@ Abstract
LiDAR point cloud segmentation is central to autonomous driving and 3D scene understanding. While voxel- and point-based methods dominate recent research due to their compatibility with deep architectures and ability to capture fine-grained geometry, they often incur high computational cost, irregular memory access, and limited runtime efficiency due to scaling issues. In contrast, range-view methods, though relatively underexplored - can leverage mature 2D semantic segmentation techniques for fast and accurate predictions. Motivated by the rapid progress in Visual Foundation Models (VFMs) for captioning, zero-shot recognition, and multimodal tasks, we investigate whether SAM2, the current state-of-the-art VFM for segmentation tasks, can serve as a strong backbone for LiDAR point cloud segmentation in the range view representations. We present RangeSAM, to our knowledge, the first range-view framework that adapts SAM2 to 3D segmentation, coupling efficient 2D feature extraction with projection/back-projection to operate on point clouds. To optimize SAM2 for range-view representations, we implement several architectural modifications to the encoder: (1) a novel Stem module that emphasizes horizontal spatial dependencies inherent in LiDAR range images, (2) a customized configuration of Hiera Blocks tailored to the geometric properties of spherical projections, and (3) an adapted Window Attention mechanism in the encoder backbone specifically designed to capture the unique spatial patterns and discontinuities present in range-view pseudo-images. Our approach achieves competitive performance on SemanticKITTI while benefiting from the speed, scalability, and deployment simplicity of 2D-centric pipelines. This work highlights the viability of VFMs as general-purpose backbones for point cloud segmentation and opens a path toward unified, foundation-model-driven LiDAR segmentation. Results lets us conclude that range-view segmentation methods using VFMs leads to promising results.
@@ 1 Introduction
Reconstructing urban scenes along with its 2D and 3D semantics is an important task in computer vision with numerous applications. In the context of creating a photorealistic simulator for autonomous systems like self-driving vehicles, drones, mobile robots, etc.—it becomes crucial to holistically represent all aspects of the scene relevant to driving. This includes tasks such as synthesizing images from interpolated and extrapolated viewpoints in real-time, reconstructing 2D and 3D semantics, generating depth information, and tracking dynamic objects. Among these perception tasks, semantic segmentation of LiDAR point clouds is fundamental, enabling real-time differentiation between vehicles, road signs, pedestrians, and infrastructure.
In recent years, research has focused on voxel- and point-based segmentation methods [57, 96, 97, 115, 34, 78]. While these approaches achieve strong performance, they impose substantial computational and memory costs on large-scale outdoor LiDAR data [63, 95, 105, 88, 83, 60, 23, 80, 61] and are challenged by the irregular, unordered nature of point clouds, which complicates efficient processing and network design. In contrast, range-view segmentation [41, 15, 3, 54, 62], projecting the 3D point cloud into a dense 2D image-like representation offers an attractive alternative. It enables reuse of mature 2D semantic segmentation models, reduces memory footprint, and supports faster inference. Despite these advantages, range-view methods have often been overlooked or deemed inferior due to earlier limitations in handling occlusions, resolution loss, and limited context modeling. Recent advances in deep learning, including attention mechanisms [82], multi-scale fusion strategies [113], and context-aware architectures [53], invite a reassessment of the range-view paradigm.
@@
  • We present the experimental results obtained using our proposed RangeSAM methodology, which demonstrates competitive performance on the SemanticKITTI benchmark [5], achieving results comparable to state-of-the-art methods in the field.
@@ 3 Methodology
This section introduces our projection-based LiDAR segmentation model leveraging VFMs - RangeSAM. We detail the range-projecting preprocessing, the modifications we contribute to the models achitecture and postprocessing procedures. The overall processing pipeline is shown in Figure 1, while Figure 23 and 4 provides a detailed schematic of the stem module, the SAM2-based encoder, and the decoder components.
@@ 3.2 Model Architecture
SAM2 represents a state-of-the-art VFM trained on the SA-V dataset, constituting the largest video segmentation dataset available at the time of its release [64]. SAM2-UNet [100] has achieved superior performance across multiple segmentation tasks, including camouflaged object detection, salient object detection, and mirror detection, while maintaining architectural simplicity. RangeSAM, adopts a similar architectural paradigm (see Figure 2), incorporating Receptive Field Blocks [46] for feature decoding to enhance segmentation performance.
@@
Table 1: Comparisons among state-of-the-art LiDAR range view semantic segmentation approaches on the test set of SemanticKITTI [5].
SqSeg[91] $\times$ 30.8 68.3 18.1 5.1 4.1 4.8 16.5 17.3 1.2 84.9 28.4 54.7 4.6 61.5 29.2 59.6 25.5 54.7 11.2 36.3
SqSegV2[93] $\times$ 39.6 82.7 21.0 22.6 14.5 15.9 20.2 24.3 2.9 88.5 42.4 65.5 18.7 73.8 41.0 68.5 36.9 58.9 12.9 41.0
RangeNet++[55] $\times$ 52.2 91.4 25.7 34.4 25.7 23.0 38.3 38.8 4.8 91.8 65.0 75.2 27.8 87.4 58.6 80.5 55.1 64.6 47.9 55.9
SqSegV3[101] $\times$ 55.9 92.5 38.7 36.5 29.6 33.0 45.6 46.2 20.1 91.7 63.4 74.8 26.4 89.0 59.4 82.0 58.7 65.4 49.6 58.9
3D-MiniNet[2] $\times$ 55.8 90.5 42.3 42.1 28.5 29.4 47.8 44.1 14.5 91.6 64.2 74.5 25.4 89.4 60.8 82.8 60.8 66.7 48.0 56.6
SalsaNext[21] $\times$ 59.5 91.9 48.3 38.6 38.9 31.9 60.2 59.0 19.4 91.7 63.7 75.8 29.1 90.2 64.2 81.8 63.6 66.5 54.3 62.1
@@
MPF[1] $\times$ 55.5 93.4 30.2 38.3 26.1 28.5 48.1 46.1 18.1 90.6 62.3 74.5 30.6 88.5 59.7 83.5 59.7 69.2 49.7 58.1
FIDNet[114] $\times$ 59.5 93.9 54.7 48.9 27.6 23.9 62.3 59.8 23.7 90.6 59.1 75.8 26.7 88.9 60.5 84.5 64.4 69.0 53.3 62.8
@@ 4 Dataset and implementation details
In this section, we describe the datasets used, the training configurations applied, and the evaluation metrics employed.
@@ 4.3 Training strategy:
All experiments are done on 8 40GB A100 GPUs. We employed a per-GPU batch size of 2, yielding an effective global batch size of 16 across the distributed setup. To stabilize training, a linear warm-up of 5 epochs is applied during the first epoch, followed by a cosine annealing schedule [28] for gradual learning-rate decay. Recent studies [41, 3, 97, 98] have demonstrated that transformer-based architectures for point cloud processing exhibit significant performance gains when leveraging pretraining and transfer learning methodologies. Building upon the training protocols established in [41, 3], we adopt an extensive pretraining strategy on the nuScenes dataset [9]. Our proposed training framework consists of the following components: To ensure dataset compatibility and enable unified training, we converted the nuScenes dataset to the SemanticKITTI format. For our experiments, we employed the SAM2-tiny backbone as the base architecture and trained the model for 60 epochs on the unified dataset. Empirical evaluation revealed that utilizing heavier SAM2 backbone variants did not yield statistically significant performance improvements, while substantially increasing computational overhead and per-frame inference latency due to the increased parameter count. This finding guided our selection of the lightweight SAM2-tiny variant to optimize the trade-off between model performance and computational efficiency. The final model architecture comprises approximately 30 million parameters, providing an optimal balance between representational capacity and computational requirements for real-time applications.
Optimization: We use the AdamW optimizer with two different schemes: 1) Hiera backbone is trained with learning rate $lr=0.0004$ and the weight decay $\alpha=0.001$, 2) The rest of the model is trained with learning rate $lr=0.001$ and the weight decay $\alpha=0.0001$
@@ 4.4 Evaluation
We present the experimental results obtained using our proposed RangeSAM methodology. Our approach demonstrates competitive performance with the SemanticKITTI benchmark (see Table 1), achieving results that are comparable to state-of-the-art methods in the field.
@@ 4.5 Quantitative evaluation
Table 2 compares our model’s IoU performance against established baseline methods. While recent state-of-the-art architectures achieve mean IoUs of $65–75\%$ our model reaches $60.9\%$. This result shows that we successfully capture the most prevalent semantic classes, even though we underperform on rarer or smaller objects. Moreover, unlike other methods that do not use any vision foundation models, our approach integrates the SAM2 model [64]—the current state of the art in semantic segmentation—demonstrating its effectiveness despite the modest mIoU gain. The quantitative results presented in Table!1 can be summarized as follows:
@@ 4.6 Qualitative evaluation
  • (b) Suburban environment: Dominant classes—cars, vegetation and fences—are correctly segmented. The model struggles with transmission towers, which were not represented in the training set, highlighting the challenge of unseen classes.
@@ 5 Conclusion
This work establishes the efficacy of VFMs for projection-base LiDAR point cloud semantic segmentation tasks. Despite the substantial domain gap between natural RGB images and range-view representations, our findings demonstrate that principled architectural adaptations tailored to the geometric properties of 3D point clouds can effectively bridge this modality shift and achieve competitive performance. Our contribution is a novel adaptation of the SAM2 architecture optimized for range-view point cloud data through targeted modifications in encoder and decoder. We develop a comprehensive training methodology that systematically integrates augmentation strategies and training protocols from prior works [15, 41, 97, 3]. Our empirical analysis reveals that extensive pretraining on 2D datasets, as proposed in [3], is unnecessary when employing visual foundation models (VFMs). Instead, we demonstrate that multi-dataset training across diverse 3D benchmarks yields superior performance gains. To facilitate reproducibility and future research, we will release both source code and model weights.
Future Work: A primary limitation of our current approach is computational complexity, particularly regarding parameter count. The convolution-free Hiera backbone [70] presents inherent challenges for point cloud segmentation due to data sparsity, which necessitates convolutional operations to introduce essential locality bias. While our experiments demonstrate that Receptive Field Blocks [46] are critical for achieving competitive performance, they constitute the primary computational bottleneck, contributing about 30 million parameters. This parameter overhead currently precludes deployment in real-time applications, representing a key area for future optimization efforts.

### v3

# RangeSAM: On the Potential of Visual Foundation Models for Range-View represented LiDAR segmentation

@@ Abstract
LiDAR point cloud segmentation is central to autonomous driving and 3D scene understanding. While voxel- and point-based methods dominate recent research due to their compatibility with deep architectures and ability to capture fine-n geometry, they often incur high computational cost, irregular memory access, and limited runtime efficiency due to scaling issues. In contrast, range-view methods, though relatively underexplored - can leverage mature 2D semantic segmentation techniques for fast and accurate predictions. Motivated by the rapid progress in Visual Foundation Models (VFMs) for captioning, zero-shot recognition, and multimodal tasks, we investigate whether SAM2, the current state-of-the-art VFM for segmentation tasks, can serve as a strong backbone for LiDAR point cloud segmentation in the range view representations. We present RangeSAM, to our knowledge, the first range-view framework that adapts SAM2 to 3D segmentation, coupling efficient 2D feature extraction with projection/back-projection to operate on point clouds. To optimize SAM2 for range-view representations, we implement several architectural modifications to the encoder: (1) a novel Stem module that emphasizes horizontal spatial dependencies inherent in LiDAR range images, (2) a customized configuration of Hiera Blocks tailored to the geometric properties of spherical projections, and (3) an adapted Window Attention mechanism in the encoder backbone specifically designed to capture the unique spatial patterns and discontinuities present in range-view pseudo-images. Our approach achieves competitive performance on SemanticKITTI while benefiting from the speed, scalability, and deployment simplicity of 2D-centric pipelines. This work highlights the viability of VFMs as general-purpose backbones for point cloud segmentation and opens a path toward unified, foundation-model-driven LiDAR segmentation. Results lets us conclude that range-view segmentation methods using VFMs leads to promising results.
@@ 1 Introduction
Reconstructing urban scenes along with its 2D and 3D semantics is an important task in computer vision with numerous applications. In the context of creating a photorealistic simulator for autonomous systems like self-driving vehicles, drones, mobile robots, etc.—it becomes crucial to holistically represent all aspects of the scene relevant to driving. This includes tasks such as synthesizing images from interpolated and extrapolated viewpoints in real-time, reconstructing 2D and 3D semantics, generating depth information, and tracking dynamic objects. Among these perception tasks, semantic segmentation of LiDAR point clouds is fundamental, enabling real-time differentiation between vehicles, road signs, pedestrians, and infrastructure. In recent years, research has focused on voxel- and point-based segmentation methods [57, 94, 95, 113, 34, 78]. While these approaches achieve strong performance, they impose substantial computational and memory costs on large-scale outdoor LiDAR data [63, 93, 103, 88, 83, 60, 23, 80, 61] and are challenged by the irregular, unordered nature of point clouds, which complicates efficient processing and network design. In contrast, range-view segmentation [41, 15, 3, 54, 62], projecting the 3D point cloud into a dense 2D image-like representation offers an attractive alternative. It enables reuse of mature 2D semantic segmentation models, reduces memory footprint, and supports faster inference. Despite these advantages, range-view methods have often been overlooked or deemed inferior due to earlier limitations in handling occlusions, resolution loss, and limited context modeling. Recent advances in deep learning, including attention mechanisms [82], multi-scale fusion strategies [111], and context-aware architectures [53], invite a reassessment of the range-view paradigm.
@@
  • We present the experimental results obtained using our proposed RangeSAM methodology, which shows promising potential on the SemanticKITTI benchmark [5], achieving competitive results that, while not state-of-the-art, validate the viability of our approach. Additionally, we conducted an ablation study to explore how various training and data augmentation strategies affect model performance.
@@ 3 Methodology
We introduce our projection-based LiDAR segmentation model leveraging VFMs - RangeSAM. We detail the range-projecting preprocessing, the modifications we contribute to the models achitecture and postprocessing procedures. The overall processing pipeline is shown in Figure 1, while Figure 23 and 4 provides a detailed schematic of the stem module, the SAM2-based encoder, and the decoder components.
@@ 3.2 Model Architecture
SAM2 represents a state-of-the-art VFM trained on the SA-V dataset, constituting the largest video segmentation dataset available at the time of its release [64]. SAM2-UNet [98] has achieved superior performance across multiple segmentation tasks, including camouflaged object detection, salient object detection, and mirror detection, while maintaining architectural simplicity. RangeSAM, adopts a similar architectural paradigm (see Figure 2), incorporating Receptive Field Blocks (RFB)[46] for feature decoding to enhance segmentation performance.
@@
Table 1: Comparisons among different Hiera backbones on the test set of SemanticKITTI [5]. Interestingly, the SAM2-tiny model outperforms its larger counterparts despite having fewer parameters, demonstrating that model capacity does not necessarily correlate with performance in this task domain.
Backbone Params(M) mIoU car bicy moto truc o.veh ped b.list m.list road park walk o.gro build fenc veg trun terr pole sign
@@
Table 2: Experiments with and without training augmentations. We observe a strong performance gain when introducing the proposed training augmentations from [41]. Comparisons among SAM2-tiny on SemanticKITTI test set [5].
Augs. mIoU car bicy moto truc o.veh ped b.list m.list road park walk o.gro build fenc veg trun terr pole sign
[…]
Table 3: Experiments with and without transfer-learning techniques using Cityscapes dataset [20]. We observe that overall our model does not benefit from extensive transfer-learning strategies introduced by [3]. Comparisons among SAM2-tiny on SemanticKITTI [5] test set.
Pretrain mIoU car bicy moto truc o.veh ped b.list m.list road park walk o.gro build fenc veg trun terr pole sign
[…]
Table 4: Comparisons among state-of-the-art LiDAR range view semantic segmentation approaches on the validation sequence of SemanticKITTI [5]. We compare our best performing model using the SAM2 with the tiny Hiera backbone.
RangeNet++[55] (IROS19) $\times$ 52.2 91.4 25.7 34.4 25.7 23.0 38.3 38.8 4.8 91.8 65.0 75.2 27.8 87.4 58.6 80.5 55.1 64.6 47.9 55.9
SqSegV3[99] (ECCV20) $\times$ 55.9 92.5 38.7 36.5 29.6 33.0 45.6 46.2 20.1 91.7 63.4 74.8 26.4 89.0 59.4 82.0 58.7 65.4 49.6 58.9
3D-MiniNet[2] (IROS20) $\times$ 55.8 90.5 42.3 42.1 28.5 29.4 47.8 44.1 14.5 91.6 64.2 74.5 25.4 89.4 60.8 82.8 60.8 66.7 48.0 56.6
SalsaNext[21] (ISVC20) $\times$ 59.5 91.9 48.3 38.6 38.9 31.9 60.2 59.0 19.4 91.7 63.7 75.8 29.1 90.2 64.2 81.8 63.6 66.5 54.3 62.1
[…]
MPF[1] (WACV21) $\times$ 55.5 93.4 30.2 38.3 26.1 28.5 48.1 46.1 18.1 90.6 62.3 74.5 30.6 88.5 59.7 83.5 59.7 69.2 49.7 58.1
FIDNet[112] (IROS21) $\times$ 59.5 93.9 54.7 48.9 27.6 23.9 62.3 59.8 23.7 90.6 59.1 75.8 26.7 88.9 60.5 84.5 64.4 69.0 53.3 62.8
@@ 4 Dataset and implementation details
We now describe the datasets used, the training configurations applied, and the evaluation metrics employed.
@@ 4.3 Training strategy:
All experiments are done on 8 40GB A100 GPUs. We employed a per-GPU batch size of 2, yielding an effective global batch size of 16 across the distributed setup. To stabilize training, a linear warm-up of 5 epochs is applied during the first epoch, followed by a cosine annealing schedule [28] for gradual learning-rate decay. Recent studies [41, 3, 95, 96] have demonstrated that transformer-based architectures for point cloud processing exhibit significant performance gains when leveraging pretraining and transfer learning methodologies. Building upon the training protocols established in [41, 3], we adopt an extensive pretraining strategy on the nuScenes dataset [9]. Our proposed training framework consists of the following components: To ensure dataset compatibility and enable unified training, we converted the nuScenes dataset to the SemanticKITTI format. For our experiments, we employed the SAM2-tiny backbone as the base architecture and trained the model for 60 epochs on the unified dataset. Empirical evaluation revealed that utilizing heavier SAM2 backbone variants did not yield statistically significant performance improvements (table 1), while substantially increasing computational overhead and per-frame inference latency due to the increased parameter count. This finding guided our selection of the lightweight SAM2-tiny variant to optimize the trade-off between model performance and computational efficiency. The final model architecture comprises approximately 63 million parameters, providing an optimal balance between representational capacity and computational requirements for real-time applications.
Optimization: We use the AdamW optimizer with two different schemes: 1) Hiera backbone is trained with learning rate $lr=0.0004$ and the weight decay $\alpha=0.001$, 2) The rest of the model is trained with learning rate $lr=0.001$ and the weight decay $\alpha=0.0001$.
@@ 4.4 Evaluation
We present the experimental results obtained using our proposed RangeSAM methodology. Table 4 presents quantitative results obtained on the validation sequence 08 of the SemanticKITTI dataset [5]. Tables 1, 2, and 3 report performance metrics evaluated on the official SemanticKITTI test set [5]. Our approach demonstrates competitive performance with the SemanticKITTI benchmark (see Table 4), achieving results that are comparable to state-of-the-art methods in the field.
@@ 4.5 Quantitative evaluation
Table 1 presents a comparative analysis of different SAM2 Hiera backbones. Due to computational constraints, our evaluation is limited to SAM2 small and tiny variants. The results demonstrate comparable mIoU performance between the Hiera tiny and small backbones, with the tiny variant offering computational advantages through reduced parameter count, resulting in faster inference times. Table 4 compares our model’s IoU performance against established baseline methods. While recent state-of-the-art architectures achieve mean IoUs of $65–75\%$ our model reaches $60.9\%$. This result shows that we successfully capture the most prevalent semantic classes, even though we underperform on rarer or smaller objects. Moreover, unlike other methods that do not use any vision foundation models, our approach integrates the SAM2 model [64], the current state of the art in semantic segmentation, demonstrating its effectiveness despite the modest mIoU gain. The quantitative results presented in Table 4 can be summarized as follows:
@@ 4.6 Qualitative evaluation
  • (b) Suburban environment: Dominant classes like cars, vegetation and fences are correctly segmented. The model struggles with transmission towers, which were not represented in the training set, highlighting the challenge of unseen classes.
@@
4.7 Ablation Study
We conduct a comprehensive ablation study to systematically evaluate the impact of different training strategies on model performance.
Transfer learning: We followed a transfer learning approach established by other baseline methods[3] in this area of research. For the experiments in Table 3, we pretrained our model on the Cityscapes [20] dataset for 25 epochs to enhance feature learning before training on the target dataset. However, this approach led to reduced performance. We believe this may be due to the SAM2 model being extensively pretrained on large-scale image datasets, which could have caused a domain mismatch affecting transfer learning. Further investigation is needed to confirm this hypothesis and explore other adaptation strategies.
Range View Augmentations: The results (see Table 2) indicate that adding augmentations introduced by [41] to the training process improves performance across nearly all datasets, achieving a 10% gain in mIoU.
@@ 5 Conclusion
This work establishes the efficacy of VFMs for projection-base LiDAR point cloud semantic segmentation tasks. To the best of our knowledge, this is the first framework to leverage VFMs for range-view projected LiDAR point cloud segmentation. Despite the substantial domain gap between natural RGB images and range-view representations, our findings demonstrate that principled architectural adaptations tailored to the geometric properties of 3D point clouds can effectively bridge this modality shift and achieve competitive performance. Our contribution is a novel adaptation of the SAM2 architecture optimized for range-view point cloud data through targeted modifications in encoder and decoder. We develop a comprehensive training methodology that systematically integrates augmentation strategies and training protocols from prior works [15, 41, 95, 3]. Our empirical analysis reveals that extensive pretraining on 2D datasets, as proposed in [3], is unnecessary when employing visual foundation models (VFMs). Instead, we demonstrate that multi-dataset training across diverse 3D benchmarks yields superior performance gains. To facilitate reproducibility and future research, we will release both source code and model weights.
Future Work: A primary limitation of our current approach is computational complexity, particularly regarding parameter count. The convolution-free hiera backbone [70] presents inherent challenges for point cloud segmentation due to data sparsity, which necessitates convolutional operations to introduce essential locality bias. While our experiments demonstrate that RFBs[46] are critical for achieving competitive performance, they constitute the primary computational bottleneck. This parameter overhead currently precludes deployment in real-time applications, representing a key area for future optimization efforts.